OpenStudy (anonymous):

Rewrite using only sinx and cosx. Sin 2x - cos 2x

OpenStudy (campbell_st):

well you need the double angle information \[\sin(2x) = 2\sin(x)\cos(x)\] and there are a number of options for cos. here is one \[\cos(2x) = \cos^2(x) - \sin^2(x)\] substitute and simplify where possible